OpenStudy (anonymous):

what is the limit as x approaches negative infinity for e^(-x^2)?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

negative infinity is still squared and becomes positive and e^(-infinity) is zero

OpenStudy (anonymous):

infinity times infinity is still infinity but since its squared it cancels out the negtive but becomes more infinty than infinity >.>

OpenStudy (anonymous):

depends on whether the equation is as written or is e^((-x)^2) In that equation, the negatives cancel out and its infinity. As you wrote it, the negative does not cancel and its 0.

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} e ^{-x^{2}} = e ^{-(-\infty)^{2}} = e^{\infty^{2}} = e^{\infty} = \infty\]